OpenStudy (anonymous):

using the chemical stoichiometry, determine the number of moles of carbonic acid that can be produced from 4 mol of NaHCO3 and 8 mol HCI

OpenStudy (aaronq):

you need to first write an equation and balance it. The use the stoichiometric coefficients to find moles produced. e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) -------------------------------------------------------- For example: if you have moles of B= 2, how many moles of C did you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_c=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) --------------------------------------------------------

OpenStudy (aaronq):

write the balanced equation, use the coefficients and a ratio (as posted above) to solve for moles of whatever species you're looking for

OpenStudy (aaronq):

idk, you need to write the balanced equation first

OpenStudy (aaronq):

you'll essentially have: \(\dfrac{n_{limiting \;reagent}}{coefficient}=\dfrac{n_{carbonic \;acid}}{coefficient}\)

OpenStudy (aaronq):

did you write the equation yet?

OpenStudy (aaronq):

it doesn't have to be, you just need to pay attention in class or read your book.

OpenStudy (anonymous):

i do but i bi-lingual nd english is hard for me so writing down takes me hrs

OpenStudy (anonymous):

@mebs using the chemical stoichiometry, determine the number of moles of carbonic acid that can be produced from 4 mol of NaHCO3 and 8 mol HCI can u help me plz

OpenStudy (aaronq):

lol

OpenStudy (anonymous):

so i know i would really appreciate your help?@aaronq

OpenStudy (aaronq):

write down the balanced equation and i'll help you further

OpenStudy (anonymous):

its okay:( i dont know how nd u dont gunna help me @aaronq thank u

OpenStudy (anonymous):

@Lena772 do u get the problems

OpenStudy (aaronq):

they gave you the formulas, all you need to do it put them in order and balance the number of atoms

OpenStudy (lena772):

Which one do you need help with, I'm confused.

OpenStudy (aaronq):

it's the same question throughout: "using the chemical stoichiometry, determine the number of moles of carbonic acid that can be produced from 4 mol of NaHCO3 and 8 mol HCI"

OpenStudy (anonymous):

@Lena772 using the chemical stoichiometry, determine the number of moles of carbonic acid that can be produced from 4 mol of NaHCO3 and 8 mol HCI this one

OpenStudy (anonymous):

i need to find the the _____ mol H2CO3 and how much excess reactant remains after the reaction ______ mol?

OpenStudy (lena772):

What topic in Chemistry is this?

OpenStudy (lena772):

Molar Ratios?

OpenStudy (anonymous):

So our balanced equation is \[NaHCO _{3} + HCl \rightarrow NaCl + H _{2}CO\] . I'm assuming this is limiting reagent question to see the minimum amount of product in this case carbonic acid we can produce

OpenStudy (anonymous):

To start we have 4 mol of NaHCO3 and 8 mol HCl . We want to know the minimum amount which could be produced and to get that we use the mole ratio that produces the least amount of the product. So in this case everything is in a 1:1 ratio. Se we learn 4 mol of NaHCO3 will give 4 mol of H2CO3 ; and 8 mol HCl will produce 8 mol H2CO3 . We want to use the least amount of mole we can gain. So 4 mol H2CO3 produced making NaHCO3 our limiting reagent

OpenStudy (anonymous):

@emanuelfarias

OpenStudy (anonymous):

@GavinxFiasco what would be the answeer?

OpenStudy (anonymous):

because i need in mol H2CO3?

OpenStudy (anonymous):

Oh i'm seeing you need to know how much excess reactant as well so we need to find how many moles of HCl which should be the same. If it's all a 1:1 mole ratio

OpenStudy (anonymous):

4.0 mol of everything it seems by the mol to mol ratio

OpenStudy (anonymous):

if we went to how many grams this would all be different

OpenStudy (anonymous):

yeah i need to know that excess reactant

OpenStudy (anonymous):

4 mol of excess reactant as well

OpenStudy (anonymous):

well it dosent want it grams does it? it says to determine the number of moles of carbonic acid produced from 4 mol of NaHCO3 and 8 Mol of HCI so i need to find it in ____mol H2CO3?

OpenStudy (anonymous):

4 mols of Carbonic acid

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and 4 mols of excess reactant

OpenStudy (anonymous):

wait what so 4 mol H2CO3 is that answer ???

OpenStudy (anonymous):

it dosent look right

OpenStudy (anonymous):

@Lena772 chapter 9 yeah mole rations can u help me

OpenStudy (anonymous):

yes 4 mols is the answer

OpenStudy (anonymous):

okay your right? now i need to find how much excess reactant remains after the reaction? _____ mol?

OpenStudy (anonymous):

@GavinxFiasco

OpenStudy (anonymous):

for the same problem?

OpenStudy (anonymous):

@GavinxFiasco

OpenStudy (anonymous):

thats the last think :)@GavinxFiasco

OpenStudy (anonymous):

I believe it's still 4 moles because you use your limiting reagent to find how many moles of your excess you will use and in this case it' s 4 mols .

OpenStudy (anonymous):

Since originally you had 8 mol then subtract the 4 you use leaves 4 mols

OpenStudy (anonymous):

@GavinxFiasco CORRECt:) yay

OpenStudy (anonymous):

You're welcome ^.^

OpenStudy (anonymous):

@GavinxFiasco ur good given the following chemical, determine how many grams of N2 are produce by 9.24g of H2O2 and 6.56g of N2H4?

OpenStudy (anonymous):

@GavinxFiasco

OpenStudy (anonymous):

Alright another limiting reagent question. I think you should open it as a new question because this thread is sort of long

OpenStudy (anonymous):

@emanuelfarias

OpenStudy (anonymous):

oh k

OpenStudy (anonymous):

idk how to that u open? yeah @GavinxFiasco

OpenStudy (anonymous):

just close this question an make a new one @emanuelfarias